SYLLABUS FOR MATH 410, FALL 2009

Monday, Wednesday, Friday 11:00 a.m. - 11:50 a.m. Classroom: MTH 0103

Prerequisites: MATH240 and MATH241, with grade of C or better; and permission of department.

Text: Advanced Calculus, 2nd Edition, by P.M. Fitzpatrick. Published by Brooks/Cole. ISBN: 0534376037 (Required)

Instructor: Wojciech Czaja Office: MTH 4406 Office Hours: Monday, Wednesday, Friday 12:00 p.m. - 12:50 p.m. or by appointment Phone: (301) 405 - 5106 Email: wojtek (at) math (dot) umd (dot) edu (Monday - Friday) Class Web Page: www.math.umd.edu/∼wojtek/410 09.htm

Exams: There will be 3 midterm exams during the semester, each for the total of 100 points. They are scheduled on October 2, October 23, and November 23. The lowest score (including absences) is dropped.

The final exam is scheduled for Wednesday, December 16, 8:00 am - 10:00 am. It is worth 200 points.

Calculators: Calculators are NOT allowed on midterm tests, nor on the final.

Homework: 7 homeworks will be given during the semester. Each is worth 20 points. The lowest 2 scores (including absences) will be dropped. Thus, the total weight of homeworks is 100 points.

Grading: The maximum point total is 600 points. The total used to calculate the final grade is the minimum of the two numbers: 600 points and the highest score in the class. The setting of letter grades will be based on the number of points and will be not worse than: 50% - D, 60% - C, 75% - B, 90% - A.

Academic integrity: The University makes me remind you about its academic integrity policies. So I do. Nobody, however, has to remind me that part of my job is to make sure that these policies are obeyed.

Attendance and absences: You are responsible for the material covered in class, whether you attend or not. You are also responsible for the announcements made during class; they may include changes in the syllabus. The instructor will adhere strictly to the official university policy on makeup exams. In particular, it is the policy of the University to excuse the absences of students that result from the following causes: illness of the student, or illness of a dependent; religious observance (where the nature of the observance prevents the student from being present during the class period); participation in university activities at the request of University authorities; and compelling circumstance beyond the student’s control. Students claiming excused absence must apply in writing and furnish documentary support for their assertion that absence resulted from one of these causes. Moreover, foreseeable absences (such as those resulting from religious holidays or participation in university-sponsored events) must be submitted in writing to the instructor by Sept. 15.

Disabilities: If you have a disability disallowing you to test under the usual time or in-class conditions, you may +contact the office of Disabled Students Services (DSS) in Shoemaker. (Please let me know by Sept. 15 if you think you may require these services.) If they assess you as meriting private conditions and/or extra time, then you may arrange to take your tests at DSS, with extra time as they indicate. You must arrange this well in advance of a test (in particular: no retakes).

Emergency closures: In case of an emergency that closes the University for an extended period of time (for example, due to inclement weather), be sure to access your email for instructions from me. Also check the University’s home page or call 301-405-SNOW for snow closure information. Should any classes or exams be cancelled, please check the class schedule page for updated schedule information.

Course evaluations: Your participation in the evaluation of courses through CourseEvalUM is important to us, and helps improve teaching and learning at the University. CourseEvalUM will be open for you to complete your evaluations for the Fall semester courses between Tuesday, December 1, and Sunday, December 13. Please provide feedback on the course, the professor, and the TA before noon on Sunday, December 13, 2009. Evaluations are anonymous and will not be available to faculty and TAs until next semester, so they cannot possibly affect your grade.